OpenStudy (anonymous):

x^2+3x-10=0 the solution(s) is/are x=?

OpenStudy (anonymous):

-5 , 2

jimthompson5910 (jim_thompson5910):

Find two numbers that multiply to -10 and add to 3. These two numbers are 5 and -2. So that gives us (x+5)(x-2)=0. Use the zero product property to then solve to get x+5=0 or x-2=0. We then get x=-5 or x=2

OpenStudy (gw2011):

Factor the equation to get: (x+5)(x-2)=0 x+5=0 x=-5 x-2=0 x=2